OpenStudy (anonymous):

Evaluate the expression: -5(18 – 7) + 9

OpenStudy (anonymous):



OpenStudy (anonymous):

Almost. -5*18 = -90 -5*-7 = +35 +9 = +9 Glue them all together and you get -90+35+9=what?